OpenStudy (anonymous):

How many grams of magnesium oxide can be produced when 97.2 g Mg react with 88.5 g O2? Unbalanced equation: Mg + O2 → MgO

OpenStudy (rogue):

First balance the equation :)

OpenStudy (anonymous):

I dont know how...

OpenStudy (rogue):

Mg + O2 → MgO Well you see that there are 2 oxygens on the reactant side while there is only 1 oxygen on the product side. To balance the oxygen, you can multiply the MgO on the product side by 2. So you have Mg + O2 → 2MgO Now if you take a look at the Magnesiums, there are 2 on the product, but only 1 on the reactant. You can multiply the magnesium on the reactant side by 2 to get both sides to equal with respect to magnesium. 2Mg + O2 → 2MgO Are there equal amounts of each atom on both sides? Yes! That is the final balanced equation.

OpenStudy (anonymous):

So 2Mg + O2 ---> 2MgO is the balanced equation?

OpenStudy (rogue):

Mhm, do you know where to go next?

OpenStudy (anonymous):

Not really I am kind of lost, this is yesterdays busywork for chem and I had the flu therefor I was not there

OpenStudy (rogue):

Oh, okay. Can you tell me what you interpret from the balanced equation?

OpenStudy (anonymous):

2Magnesium + Oxygen 2 ----> 2MagnesiumOxygen

OpenStudy (anonymous):

Hello? Look I really would just appriciate it explained to me in detail I have to go to bed soon and I am kind of in a time crunch here

OpenStudy (rogue):

Yes, but what does it say about the actual reaction that takes place? When 2 moles of magnesium reacts with 1 mole of oxygen gas, 2 moles of magnesium oxide will be formed. So let's say you mix 10 moles of Mg with 5 moles, you'll get 10 moles of MgO2. You can solve your problem by converting the masses given to moles and using this ratio.

OpenStudy (rogue):

Sorry for the late replies, my printer got in a jam and I'm a bit busy atm.

OpenStudy (anonymous):

Ok well I am confused

OpenStudy (rogue):

How many moles of Mg are in 97.2 g of Mg? How many moles of O2 are in 88.5 g of O2?

OpenStudy (anonymous):

@dumbcow please help!

OpenStudy (rogue):

I see how it is! It's okay... ;)

OpenStudy (anonymous):

I just am not understanding you! I am sorry!

OpenStudy (anonymous):

@ganeshie8 please help!

OpenStudy (anonymous):

@AccessDenied please help!

OpenStudy (anonymous):

@Outkast3r09 please help!! ANYBODY!

OpenStudy (anonymous):

@surdawi please help!

OpenStudy (anonymous):

@Rogue I NEED HELP!

OpenStudy (rogue):

-_- sigh...

OpenStudy (anonymous):

?

OpenStudy (anonymous):

How many grams of calcium phosphate can be produced when 89.3 grams of calcium chloride reacts with excess sodium phosphate? Unbalanced equation: CaCl2 + Na3PO4 → NaCl + Ca3(PO4)2 Show, or explain, all of your work along with the final answer. balanced equation 3 CaCl2 + 2 Na3PO4 = 6 NaCl + Ca3(PO4)2 moles CaCl2 = 89.3 g /110.984 g/mol=0.805 moles Ca3(PO4)2 = 0.805/3 =0.268 mass Ca3(PO4)2 = 0.268 mol x 310.17 g/mol=83.1 g is this at least right

OpenStudy (anonymous):

I am going to cry

OpenStudy (rogue):

How many grams of magnesium oxide can be produced when 97.2 g Mg react with 88.5 g O2? 2Mg + O2 --> 2MgO The atomic mass of magnesium is ~24.31 grams/mole (1 mole of magnesium weighs 24.31 grams). So in 97.2 grams of magnesium, there is 97.2/24.31 = 4 moles of magnesium. Looking at the reaction, 2 moles of magnesium yields 2 moles of magnesium oxide. So 4 moles of magnesium should yield 4 moles of magnesium oxide. You also have to consider the amount of oxygen you have. 1 mole of O2 weighs ~32 grams and so in 88.5 grams of oxygen, there are 88.5/32 = 2.77 moles. From the reaction, 1 mole of O2 yields 2 moles of MgO. So 2.77 moles of O2 should yield 5.54 moles of MgO. However, both the Mg and the O2 are reacting and there is only a limited amount of each. The Mg will run out before the oxygen and so only 4 moles of magnesium oxide can be produced. 1 mole of MgO weighs ~ 40.3 grams. So 4 moles of MgO, your final product, should weigh ~ 161.2 grams.

OpenStudy (rogue):

Does that make sense?

OpenStudy (rogue):

Yeah, you have the calcium phosphate problem right! :) It's actually a bit easier. Good job though :D

OpenStudy (anonymous):

Our equation is 2Mg + O2=2MgO also 4Mg + 2O2=4MgO 2*(24.3+16) 4*(24.3+16) 161.2 gram of MgO formed Is this right

OpenStudy (anonymous):

@Rogue ?

OpenStudy (rogue):

Yeah, that's correct.

OpenStudy (anonymous):

Thank you!!

OpenStudy (anonymous):

@Rogue are you sure?

OpenStudy (rogue):

Fairly certain :)

OpenStudy (anonymous):

@Rogue is this right as well? What is the limiting reactant when 8.4 moles of lithium react with 4.6 moles of oxygen gas? Unbalanced equation: Li + O2 → Li2O 4Li + O2 --> 2Li2O 8.4(2/2) = 8.4 4.6 O2(2/1 O2) = 9.2 g

OpenStudy (rogue):

4 moles of lithium and 1 mole of O2 would yield 2 moles of Li2O. 8.4 moles of lithium and x moles of O2 would yield 4.2 (8.4/2) moles of Li2O. x = 8.4/4 = 2.1 moles of O2 required to react with all the Lithium. We have 4.6 moles of O2, more than enough. x moles of lithium and 4.6 moles of O2 would yield 9.2 (4.6*2) moles of Li2O. x = 4*4.6 = 18.4 moles of lithium is required to completely exhaust the 4.6 moles of O2. We only have 8.4 moles of Lithium, so there isn't enough Li. Lithium is the limiting reactant.

OpenStudy (anonymous):

No.. This is not a practice exam. This is a module quiz from FLVS so using this answer will be an academic integrity violation. Plus it has a obvious mistake in it. Mr. Carlyle FLVS Chemistry Instructor This question is a violation of the OpenStudy Guidelines. The question is from an online school plus a violation by solicitation for members to unknowingly assist them cheating on an exam. 30653065chem30653065